OpenStudy (anonymous):

Find a number that is between 7/11 and __ 0.75

OpenStudy (anonymous):

A. 76/99 B. 23/33 C. 77/99 D. 25/44

OpenStudy (anonymous):

A and C are too large, D is too small

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

\[\frac{ 7 }{ 11 }=0.64\] \[\frac{ 76 }{99 }=0.77\] \[\frac{ 23 }{33 }=0.696\] \[\frac{ 25 }{44 }=0.568\] \[no. \between \frac{ 7 }{11 }and 0.75 is \frac{ 23 }{ 33 }\]